OpenStudy (anonymous):

The ratio of an object's weight on Earth to its weight on Neptune is 5 : 7. How much would a person who weighs 150 pounds on Earth weigh on Neptune?

OpenStudy (anonymous):

5/7 = 150/x x=150*7/5 = 210

OpenStudy (campbell_st):

the ratios are \[\frac{5}{7} = \frac{150}{x}\] or \[x = \frac{7 \times150}{5}\] just finish the calculation

OpenStudy (anonymous):

thank u everyone that help with my question